the sum of the numerator& denominator of a fraction is 17. If 3 is added to the numerator, the value of fraction will be 1. What is the fraction?
Found 2 solutions by solver91311, richwmiller:
Answer by solver91311(24713)



Let represent the numerator. Then let represent the denominator.

We are given that:



In order for the value of a fraction to be 1, the denominator and the numerator must be equal, therefore we now know that:



Substitute and solve:



You should be able to take it from here.

John


Answer by richwmiller(17219)

a+b=17
b=17-a
(a+3)/b=1
(a+3)/(17-a)=1
a+3=17-a
2a=14
a=7
b=10
(7+3)/10=1